# Joint State and Noise Covariance Estimation

### v1

@@ VIII Conclusion
This work presented a novel and rigorous framework for joint estimation of primary parameters and noise covariance matrices in SLAM and related computer vision estimation problems. We derived analytical expressions for (conditionally) optimal noise covariance matrix (under ML and MAP criteria) under various structural constraints on the true covariance matrix. Building on these solutions, we proposed two types of algorithms for finding the optimal estimates and theoretically analyzed their convergence properties. Our results and algorithms are quite general and can be readily applied to a broad range of estimation problems across various engineering disciplines. Our algorithms were validated through extensive experiments using linear measurement models and PGO problems.
@@ Appendix B Proof of Proposition 1
The objective function is strictly convex because log-determinant is strictly concave over $\mathbb{S}_{\succ 0}$ and the second term is linear in $P$. The equality (diagonal) and “inequality” (eigenvalue) constraints are also linear and convex, respectively. Since the objective is strictly convex, there is at most one optimal solution (i.e., if a minimizer exists, it is unique).
@@ Appendix C Proof of Theorem 1
For a given $x$, the Lagrangian of the inner subproblem in Problem 1 is given by
$\mathcal{L}(P,Q)=-\log\det P+\langle M(x),P\rangle+\langle Q,P\rangle,$ (39)
where $Q$ is the Lagrange multiplier corresponding to the semidefinite cone constraints. Observe that $\nabla_{P}\mathcal{L}(P^{\star}(x))=Q^{\star}$ for $P^{\star}(x)=M(x)^{-1}$. Since, $P^{\star}(x)$ in this case is primal feasible then optimal dual variable would $Q^{\star}=0$. Thus, $P^{\star}(x)$ is the optimal solution to the problem.
@@
$\mathcal{L}(\lambda_{1},\dots,\lambda_{m},q_{1},\dots,q_{m})=\sum_{i=1}^{m}\left[-\log\lambda_{i}+\lambda_{i}M_{ii}(x)+\lambda_{i}q_{i}\right],$ (41)
@@
$\underline{\Lambda}_{ii}(x)=\begin{cases}0&D_{ii}(x)\in[0,\lambda_{\text{max}}],\\ D_{ii}(x)-\lambda_{\max}&D_{ii}(x)\in(\lambda_{\text{max}},\infty),\\ \end{cases}$ (51)
@@
$\overline{\Lambda}_{ii}(x)=\begin{cases}\lambda_{\min}-D_{ii}(x)&D_{ii}(x)\in[0,\lambda_{\text{min}}],\\ 0&D_{ii}(x)\in(\lambda_{\text{min}},\infty).\\ \end{cases}$ (52)
It can be observed that the choice of $P^{\star}(x)$ as in (14) along with $\overline{Q}^{\star}(x)$ and $\underline{Q}^{\star}(x)$ defined above satisfy the KKT conditions and are the primal-dual optimal solutions to the porblem. This completes the proof.
@@ Appendix D Proof of Theorem 2
The proof follows from specializing the proof of [20, Theorem 4] to the problem considered in this paper. Specifically, the compactness of manifolds along with Assumptions 1 and 2 lead to the existence of of $\tilde{L}$ as described in Lemma 1. Then following the steps of the proof of [20, Theorem 4] and setting $b=2$ completes the proof.
@@ Appendix E Background Information for Section V-B
Let $\mathcal{Y}$ be a smooth submanifold in Euclidean space. Each $y\in\mathcal{Y}$ is associated with a linear subspace, called the tangent space, denoted by $T_{y}\mathcal{Y}$. Informally, the tangent space contains all directions in which one can tangentially pass through $y$. For a formal definition see [1, Definition 3.5.1, p. 34]. A smooth function $F:\mathcal{Y}\rightarrow\mathbb{R}$ is associated with the gradient $\nabla F(y)$ of $F$, as well as the orthogonal projection of $\nabla F(y)$ onto the tangent space $T_{y}\mathcal{Y}$, known as the Riemannian gradient of $F$ at $y$, denoted by $\mathrm{grad}\,F(y)$. See [1, p. 46] for more detail. The manifold $\mathcal{Y}$ is also associated with a map $\mathrm{Retr}_{y}:T_{y}\mathcal{Y}\rightarrow\mathcal{Y}$, called a retraction that enables moving from $y$ along the direction $v\in T_{y}\mathcal{Y}$ while remaining on the manifold – see [4, Definition 3.47] and [1, Definition 4.1.1] for formal definitions.
A sufficient condition for Assumption 2 to hold is for the manifolds to be compact subsets of the Euclidean space. Under this assumption we have the following lemma from [20].

### v3

@@ VIII Conclusion
This work presented a novel and rigorous framework for joint estimation of primary parameters and noise covariance matrices in SLAM and related computer vision estimation problems. We derived analytical expressions for (conditionally) optimal noise covariance matrix (under ML and MAP criteria) under various structural constraints on the true covariance matrix. Building on these solutions, we proposed two types of algorithms for finding the optimal estimates and theoretically analyzed their convergence properties. Our results and algorithms are quite general and can be readily applied to a broad range of estimation problems across various engineering disciplines. Our algorithms were validated through extensive experiments using linear measurement models and PGO problems. The results show that the state and the noise covariance matrix can be jointly estimated from the measurements (and, optionally, a prior on the covariance) with negligible computational overhead relative to standard solvers.
@@ Appendix A Wishart Prior
Remark 7 (Wishart Parameters).
It is worth noting that Algorithm 1 represents just one simple approach to setting the prior parameters. Alternative strategies can also be employed, such as learning the parameters from an offline calibration dataset, or directly specifying the value of $\nu$ independently of $k$ (with larger values of $\nu$ indicating greater confidence in the prior estimate $\Sigma_{0}$). By explicitly setting a constant value for $w_{\text{prior}}$, [^3] That is, setting the prior parameter $\nu$ as a function of $k$ in (36). the relative influence of the prior versus the measurements remains fixed. We adopted this strategy in our experiments for simplicity. However, it is important to recognize that setting $\nu$ as a function of $k$ violates the Bayesian principle that the prior should be specified before observing any data. Consequently, when $w_{\text{prior}}$ is kept fixed, unlike in standard MAP estimation the measurements will never completely dominate the prior as $k$ increases.
@@ Appendix B Proof of Proposition 1
The objective function is strictly convex because log-determinant is strictly concave over $\mathbb{S}_{\succ 0}$ and the second term is linear in PP. The equality (diagonal) and linear matrix inequality (eigenvalue) constraints are also linear and convex, respectively. Since the objective is strictly convex, there is at most one optimal solution (i.e., if a minimizer exists, it is unique).
@@ Appendix C Proof of Theorem 1
In the following, we provide primal-dual pairs that satisfy the KKT conditions for each variant. Since the problem is differentiable and convex, these pairs must be primal-dual optimal. Moreover, these solutions are unique because the objective function is strictly convex.
@@
For a given $x\in\mathcal{M}$, the Lagrangian of the inner subproblem in Problem 1 is given by
$\mathcal{L}(P,Q)=-\log\det P+\langle M(x),P\rangle-\langle Q,P\rangle,$ (44)
where $Q\succeq 0$ is the Lagrange multiplier corresponding to the semidefinite cone constraint. Recall that $M(x)\succ 0$ by construction. Verify that $P^{\star}(x)=M(x)^{-1}$ and $Q=0$ trivially satisfy the KKT conditions, and thus $P^{\star}(x)=M(x)^{-1}$ is the optimal solution to the primal problem.
@@
$\mathcal{L}(\lambda_{1},\dots,\lambda_{m},q_{1},\dots,q_{m})=\sum_{i=1}^{m}\left[-\log\lambda_{i}+\lambda_{i}M_{ii}(x)-\lambda_{i}q_{i}\right],$ (46)
@@
[…]
@@
[…]
It can be observed that the choice of $P^{\star}(x)$ as in (18) along with $\overline{Q}^{\star}(x)$ and $\underline{Q}^{\star}(x)$ defined above satisfy the KKT conditions and are the primal-dual optimal solutions to the problem. This completes the proof.
@@ Appendix D Proof of Theorem 2
Consider a strictly feasible pair $(x_{0},P_{0})$ where $x_{0}\in\mathcal{M}$ and $P_{0}\succ 0$. Let $u_{0}\in\mathbb{R}^{m}$ be any vector in the nullspace of $S(x_{0})$. For any $c>0$, the pair $(x_{0},P_{0}+c\,u_{0}u_{0}^{\top})$ remains strictly feasible. Observe that
$F_{\text{ML}}(x_{0}$ $,P_{0}+c\,u_{0}u_{0}^{\top})=$
$-\log\det\left(P_{0}+c\,u_{0}u_{0}^{\top}\right)+\langle S(x_{0}),P_{0}+c\,u_{0}u_{0}^{\top}\rangle$ (58)
$=-\log\det\left(P_{0}+c\,u_{0}u_{0}^{\top}\right)+\langle S(x_{0}),P_{0}\rangle$ (59)
$=-\log\left(1+c\,u_{0}^{\top}P_{0}^{-1}u_{0}\right)-\log\det P_{0}+\langle S(x_{0}),P_{0}\rangle$ (60)
$=-\log\left(1+c\,u_{0}^{\top}P_{0}^{-1}u_{0}\right)+F_{\text{ML}}(x_{0},P_{0}),$ (61)
where we used the matrix determinant lemma in (60). Since $P_{0}\succ 0$, we have that $u_{0}^{\top}P_{0}^{-1}u_{0}>0$. Hence, as $c\to\infty$, we have
$F_{\text{ML}}(x_{0},P_{0}+c\,u_{0}u_{0}^{\top})\to-\infty.$
Therefore, the problem is unbounded below and does not admit a solution.
@@ Appendix E Background Information for Section V-B
Let $\mathcal{Y}$ be a smooth submanifold in Euclidean space. Each $y\in\mathcal{Y}$ is associated with a linear subspace, called the tangent space, denoted by $T_{y}\mathcal{Y}$. Informally, the tangent space contains all directions in which one can tangentially pass through $y$. For a formal definition see [2, Definition 3.5.1, p. 34]. A smooth function $F:\mathcal{Y}\rightarrow\mathbb{R}$ is associated with the (Euclidean) gradient $\nabla F(y)$ of $F$, as well as the orthogonal projection of $\nabla F(y)$ onto the tangent space $T_{y}\mathcal{Y}$, known as the Riemannian gradient of $F$ at $y$, denoted by $\mathrm{grad}\,F(y)$. See [2, p. 46] for more detail. The manifold $\mathcal{Y}$ is also associated with a map $\mathrm{Retr}_{y}:T_{y}\mathcal{Y}\rightarrow\mathcal{Y}$, called a retraction that enables moving from $y$ along the direction $v\in T_{y}\mathcal{Y}$ while remaining on the manifold – see [7, Definition 3.47] and [2, Definition 4.1.1] for formal definitions.
A sufficient condition for Assumption 2 to hold is for the manifolds to be compact subsets of the Euclidean space. Under this assumption we have the following lemma [25, Lemma 1].
@@
Appendix F Proof of Theorem 3
The proof follows from specializing the proof of [25, Theorem 4] to the problem considered in this paper. Specifically, the compactness of manifolds along with Assumptions 1 and 2 lead to the existence of of $\widetilde{L}$ as described in Lemma 1. Then following the steps of the proof of [25, Theorem 4] and setting $b=2$ completes the proof.